OpenStudy (anonymous):

How do you simplify this? (3x^-1y^-2)^-2(x^2y^5)^-1 The negative exponents out of the parenthesis throw me off

OpenStudy (anonymous):

then multiply each exponent inside the parentheses with the exponent outside the parentheses, maybe that will make it easier

OpenStudy (anonymous):

\[(3x^{-1}y^{-2})^{-2}(x^2y^5)^{-1}=(3^{-2}x^2y^4)(x^{-2}y^{-5})\]

OpenStudy (anonymous):

So can i simplify this anymore? (x^2y^4/9)(1/x^2y^5)

OpenStudy (anonymous):

yes ignore the parentheses in the first line i wrote and add the exponents the parentheses are not necessary when you are multiplying

OpenStudy (anonymous):

i means add the exponents for the like terms. for example \(x^2\times x^{-2}=x^{2-2}=x^0=1\)

OpenStudy (anonymous):

So is the answer (x^4y^9/9) how do i know which goes in numerator/denominator

OpenStudy (anonymous):

That's not the answer. \[(3x^{-1}y^{-2})^{-2}(x^2y^5)^{-1}\]\[=(3^{-2}x^{(-1)\times (-2)}y^{(-2)\times (-2)})(x^{2\times (-1)}y^{5\times (-1)})\]\[=(3^{-2}x^{2}y^{4})(x^{-2}y^{-5})\]\[=3^{-2}x^{2}x^{-2}y^{4}y^{-5}\]\[=3^{-2}x^{2+(-2)}y^{4+(-5)}\]\[=...?\]

OpenStudy (anonymous):

1/9y ??

OpenStudy (anonymous):

yes